# User Location Disclosure Fails to Deter Overseas Criticism but Amplifies Regional Divisions on Chinese Social Media

### v2

@@ 2.1 High-frequency Weibo Data
We conducted real-time monitoring of each account’s timeline (i.e., their posts), each post’s first-page comment section (displaying up to the top 20 comments), and associated engagement metrics at both the post and comment levels. We focus on the first-page comments because they were instantly available to readers when a post is clicked and draw the most user attention and interaction. As shown in the top-left panel of Figure 1(a), our system scanned all monitored accounts, as well as recorded posts and comments, every five minutes. This allowed us to collect longitudinal data at both the account and post levels. When a new post appeared, the system initiated tracking and began recording data on the post and its top comments. Snapshots were captured every five minutes for the first 24 hours, and once every day for the following ten days. While most engagement occurred within the first few hours (see Figure S4 in the SI), the extended tracking window enabled us to capture delayed or atypical activity. Figure 1(b) illustrates the variables collected through scraping, including each post’s content, publication time, and engagement metrics (number of reposts, likes, and comments), as well as the first-page of the comment section (top 20 comments). For each comment and their replies, we also recorded the associated IP location tag. Figure 1(c) and Figure 1(d) show the numbers of new posts, and new comments collected within a 48-hour window, respectively. Figures 1(e) and 1(f) show that, on average, comments appearing on and after the second page receive minimal engagement.
2.2 Timeline and Location Tags
@@
The location tags reflect users’ IP addresses at the time of commenting, although we do not observe or record the IPs themselves. Overseas users are labeled by country or region, while domestic users are tagged at the provincial level (e.g., Shanghai, Jiangsu, Shaanxi). [^2] A potential concern is that users may mask their IP addresses using VPNs. in the SI, we show that fewer than 1% of users changed location tags during the study period. VPN usage among domestic users is rare and typically serves to switch between domestic and foreign IPs, not between provinces. To our knowledge, commercially available VPN services in China do not support intra-national IP masking, as all provinces operate within the same national firewall. These patterns suggest that provincial-level location tags are generally reliable. Excluding users whose location tags have changed does not alter on our findings.
@@
Post content: positive, negative, or neutral event Post LLM-assisted annotation + human validation 0.7210.7210.721
@@
Comment stance: critical of the post or not Comment Supervised BERT classifier + human validation 0.7830.7830.783
Comment receiving regionally discriminatory replies Comment LLM-assisted annotation + human validation 0.8350.8350.835
2.3 Key variables and Measurement
@@
We also classify post types based on their content. International posts are those that mention one of 123 countries or regions outside China. We construct the list of 123 countries and regions (see Table S4 in the SI) by extracting all foreign country and region names that appeared in Weibo’s post-treatment IP location tags. Local posts reference one of China’s 31 provinces. We focus on provincial references, rather than prefectures or counties, because IP location tags are reported at the country level for overseas users and the provincial level for domestic users. Posts are further categorized as depicting a positive, negative, or neutral event (from the perspective of the Chinese government) using a large language model (LLM), followed by human validation.
@@
To detect regional discrimination in replies, we use LLM-based annotation, validated by human coders, to identify whether a reply contains discriminatory content based on the commenter’s location. We also track the frequency with which a user’s IP location is mentioned in replies as an indirect measure of the salience of geographic identity in the discourse. Table 1 summarizes these variables and their corresponding measurement strategies. Table S3 in the SI provides the summary statistics.
2.4 Protecting Privacy
All data used in this study are publicly available. Nevertheless, we take additional steps to protect the privacy of commenters, many of whom are likely ordinary citizens in China. First, we do not collect any personally identifiable information, such as commenters’ IP addresses. The location tags we use are limited to the provincial level and are too coarse to identify individuals. Second, we mask all user identifiers—including usernames, avatars, and URLs—by replacing them with randomly assigned IDs to ensure complete anonymity. Third, all data are stored on a secure server at the corresponding author’s institution, with access restricted to authorized personnel only (see Table S2 in the SI for details.)
3 Identification Strategy
@@
Because an interrupted time series design is analogous to a regression discontinuity design with time as the forcing variable (Hausman and Rapson, 2018), we estimate $\tau_{\text{SAITT}}$ using local linear regression on either side of $c$, clustering standard errors at the account level. Bandwidths are selected using both the MSE-optimal procedure in rdrobust (Calonico et al., 2014) and cross-validation. We show in the SI (Figures S12 to S17) that our results are robust to a wide range of bandwidth choices.
4 Findings
Using the data and identification strategy described above, we examine two questions: (1) Did revealing user IP locations deter overseas and domestic Weibo users from participating in online discussions? (2) How did this policy shape intergroup discussions?
The policy was designed to deter participation by overseas users with “malicious intent,” who were believed to interfere in domestic discourse by spreading misinformation and expressing views that challenge the official narrative. The assumption was that once users’ IP locations became public, domestic users would be better equipped to judge the credibility of online content, thereby reducing the effectiveness of malicious foreign propaganda. At the same time, overseas users might reduce their engagement out of concern about being discredited or attacked as part of a “malicious foreign force” (see SI). To evaluate whether the policy has achieved its goal, we first distinguish between posts related to international affairs and those that are not. Posts by government and media accounts on international affairs are defined as those that explicitly mention at least one foreign location from the aforementioned country list; all other posts are classified as non-international, which include both posts covering domestic issues and general content without a clear reference to a foreign country.
(a) #Comments by Oversea Users
(b) #Comments by Domestic Users
(c) #Comments by Domestic Users on Non-International issues
(d) Proportion of Local Posts Covering Positive & Negative Events
(e) #Comments by Domestic Users:
By Commenters Origin
(f) #Critical Comments by Domestic Users:
By Commenters Origin
Figure 2: Engagement (Comments and Replies) before and after User Location Disclosure. Figures 2(a) and 2(b) show smoothed trends in the number of first-page comments from overseas and domestic users, respectively, on posts about international and non-international topics. Comment counts are aggregated over the first 48 hours after posting, and the vertical red line marks the implementation of the IP location disclosure policy at noon on April 28, 2022. Figure 2(a) shows no decline in overseas engagement. Comments on international topics rose temporarily before returning to baseline, while comments on non-international topics remained stable. By contrast, Figure 2(b) shows a sharp and sustained drop in domestic comments on non-international posts, with no change on international topics. Figure 2(c) confirms that this decline is driven by posts about local issues—defined as those mentioning a province name—while engagement with non-local topics remains stable. Figure 2(d) shows no significant change in the distribution of local post content by sentiment, ruling out shifts in post type as an explanation. Figures 2(e) and 2(f) further show that the drop in engagement is driven primarily by out-of-province users and is accompanied by a reduction in dissenting comments. In-province engagement also declined but rebounded more quickly. These patterns suggest that user location disclosure discouraged nonlocal participation in local discussions. Confidence intervals are constructed based on cluster-robust standard errors clustered at the account level.
Figure 2(a) directly tests this expectation. The blue and red smoothed lines show the average number of comments from overseas users (based on platform-identified IP locations) on international and non-international posts, respectively, before and after the policy took effect. As explained earlier, these IP locations are observable because, during the first 12 hours following the policy’s rollout, Weibo retroactively displayed location tags on historical comments. This implementation quirk allows us to classify user origin even for comments made prior to the policy change. Contrary to expectations, there is no evidence of reduced participation by overseas users. Instead, we observe a sharp increase in comments by overseas users on international topics, rising from 0.41 to 0.78 average per post immediately after the policy’s implementation, followed by a gradual return to baseline over the next ten days. This temporary surge suggests a reactive backlash rather than retreat.
One alternative explanation is that domestic users, masking their locations via Virtual Private Networks (VPNs), contributed to this surge. Yet this is unlikely: we find that IP location tags remained unchanged for 99.1% of users before and after the policy was implemented, indicating minimal switching. Furthermore, analysis in Figure S25 in the SI shows that overseas users continued to engage with sensitive topics such as the Russian invasion of Ukraine, despite facing increased attacks from domestic commenters, likely provoked by the newly visible location tags following the policy implementation.
Moreover, we find no decline in comments on non-international posts (red line in Figure 2(a)) by international users, which further weakens the deterrence hypothesis. The red line remains consistently lower than the blue line, but has much narrower confidence intervals. This is because overseas users are more likely to comment on posts related to international affairs; however, such posts are less frequently made by propaganda accounts compared to non-international posts, which explains the wider confidence intervals around the blue line.
Given that user location disclosure did not reduce participation from overseas users—whether in international or domestic discussions—the policy appears to have failed in achieving its goal of discouraging overseas engagement on topics deemed sensitive. This naturally raises the following question: if the intended targets were unaffected, what impact did the policy have on domestic users?
Figure 2(b) shows that domestic engagement declined. As in Figure 2(a), the blue and red smoothed lines represent the average number of comments on international and non-international posts, respectively—this time from domestic users, based on their platform-identified user locations. The volume of comments on international posts by domestic users remains stable following the policy’s introduction. In contrast, the average number of comments on non-international posts drops sharply (by 50%) immediately after the policy took effect. Again, the wider confidence intervals around the blue line reflect the fact that government and media accounts post far fewer international than non-international topics, resulting in greater variability in the estimated trend.
In Figure 2(c), we further examine the change in comment volume by domestic users on posts related to non-international issues, where a sharp decline followed the introduction of the policy. We divide these posts into two categories: those concerning local issues (purple lines)—defined as posts that mention 31 province names in mainland China (see Table S5 in the SI for the list)—and those addressing other, non-local topics (red lines). We define local issues in this manner because the IP location tags were assigned at the provincial level. The results show that the overall decline in the volume of comments on non-international content in Figure 2(b) is primarily driven by a drop in engagement with posts about local issues. In fact, the average number of comments per post on local issues by domestic users dropped sharply from 8.21 to approximately 0.42 after the policy was implemented. [^3] Consistent with this finding, Figure S27 in the SI shows that aggregate engagement metrics, including the total number of comments, likes, and reports, on posts about local issues also declined following the policy. However, using the aggregate metrics, we cannot distinguish whether this decrease is driven by domestic or international users. In contrast, the volume of comments on non-local issues (posts without mentioning a province name) remains relatively stable before and after the policy’s implementation.
4.2 The drop in comments on posts about local issues was not driven by changes in post content, but by reduced engagement from out-of-province commenters
We now turn to posts concerning local affairs. This is where we observe the sharpest decline in user engagement following the implementation of the location disclosure policy. A natural explanation is that the policy produced a chilling effect: by revealing users’ provincial origins, it may have discouraged them from commenting on sensitive local issues, especially those involving negative events such as accidents, scandals, or misconduct by local government officials.
To assess this possibility, we examine whether there was any change in the content of the posts themselves. If, for instance, government or media accounts began posting fewer negative or politically sensitive stories after the policy was introduced, that might explain the decline in comments. However, Figure 2(d) shows that the proportion of local posts covering positive, neutral, and negative events remained stable before and after the policy’s implementation. These negative events include major accidents, scandals, and COVID-lockdown. In fact, we show in Figure S6 in the SI that comments declined across posts featuring both positive and negative events. This suggests that the drop in engagement was not driven by changes in what was being posted, but by changes in who chose to respond.
If the declining volume of domestic commentators on posts related to local events is not driven by changes in the content of the posts themselves, what then can explain this sharp drop? We suspect that the change in location disclosure has altered the incentive for commenters discussing local affairs.
To investigate this possibility, we examine whether the effect of the policy varied depending on the commenter’s location relative to the content of the post. Specifically, we distinguish between “in-province” users—those commenting on posts about their own province—and “out-of-province” users—those commenting on posts about other regions. Figure 2(e) reveals that the volume of comments from out-of-province users declined from on average 6.10 per post to 0.42 per post, and this decline persisted for days after the policy took effect. In-province users also reduced their participation, from 1.85 per post to 0.15 per post, and their engagement rebounded quickly. This pattern indicates that the overall drop in discussion on local affairs was primarily driven by the withdrawal of commenters from outside the region.
The policy not only reduced how many people commented, but also altered the nature of the comments themselves. Figure 2(f) shows that the proportion of critical or dissenting comments on local posts fell significantly after the IP location tags were introduced. This decline was largely driven by the exit of out-of-province commenters, who had previously made up a sizable share of those expressing criticism.
In sum, while the policy had little effect on overseas users (if anything, a backlash), it substantially dampened domestic discussion of local issues. The sharpest declines came from out-of-province users, who—once their location was visible—appeared less willing to voice criticism about regions to which they did not reside. These findings raise further questions: Why did location tagging silence domestic critics more effectively than it deterred overseas participants? And how did the visibility of regional identity reshape the norms and risks of online expression within China? To answer these, we turn next to the role of regional divisions in shaping online discourse.
(a) Intensified Regional Discrimination
(b) Less Comments with Reply
(c) Reduced Coefficient of Variation
To explain why domestic users—particularly those commenting on regions other than their own—became less engaged in local discussions following the introduction of IP location disclosure, we consider three possible mechanisms.
The first, and perhaps most intuitive, explanation is a direct chilling effect. The presence of a location tag may have reminded users that their activity was being monitored by the state, thereby discouraging critical commentary. However, this explanation is unlikely to be the primary driver. As shown in Figures 2(b) (blue line) and 2(c) (red line), domestic users did not reduce their engagement on international and national (non-local) issues, including posts that attracted criticism.
A second possibility is that location tags undermined the perceived credibility of out-of-province commenters by signaling their lack of local knowledge. This interpretation aligns with the policy’s stated goal of curbing misinformation—perhaps charitably understood as discouraging speech from those unfamiliar with local context. If this mechanism were driving the observed decline, however, we would expect similar reductions in two cases: (a) overseas users commenting on domestic topics, and (b) domestic users commenting on international topics, as both groups are equally or more likely to lack firsthand knowledge. Yet we observe no such decline (Figure 2(a), red line; Figure 2(b), blue line), suggesting that concerns about credibility alone cannot explain the pattern.
A third possibility, supported by widespread anecdotal evidence, is that domestic users became more hesitant to comment on out-of-province issues due to heightened risk of backlash from local users. By making users’ provincial origins visible, the policy increased the salience of regional identity and turned comment sections into sites of intergroup tension. This mechanism should not affect comments on posts about international or national (nonlocal) issues, where location tags are less likely to provoke region-specific pushback.
Before location disclosure, critical remarks on local issues often appeared anonymous or broadly directed. Afterward, they became traceable to specific locations. Even generic observations could now be read as targeted attacks from outsiders. Figure 29(b) in the SI illustrates this dynamic: a user from Fujian, commenting on population decline in the northeast, draws on the stereotype that “mafia is everywhere” in Heilongjiang, Liaoning, and Jilin. The visible location tag triggers retaliatory responses that invoke tropes about fraud and scams in Fujian. What might have been a passing comment escalates into an interprovincial confrontation.
To assess whether this pattern was isolated or systemic, we use an LLM to detect regionally discriminatory language in second-tier replies. Figure 3(a) shows a sharp rise in such language following the introduction of location tags. As shown in Figure S26 in the SI, this increase is almost entirely driven by cross-provincial interactions; within-province exchanges remain stable. These findings suggest that the policy deepened existing regional divisions and raised the social and reputational costs of engaging in discussions about other regions.
This dynamic helps explain the observed decline in comment activity on local posts. As regional identity became more visible, users faced greater risks of backlash—not just for commenting critically, but for engaging at all. Even neutral remarks could provoke confrontation, discouraging participation and reducing the likelihood that comments would receive replies. Figure 3(b) captures this shift: after location disclosure, a smaller share of comments attracted responses, indicating thinner, more cautious engagement.
We further quantify this decline by measuring the rotation among first-page comments using the coefficient of variation (CoV) of comment floor numbers—a proxy for dynamism in the comment section. Figure 3(c) shows a clear drop in CoV following the policy rollout. As users became more hesitant to comment, especially on posts involving other provinces, top comments were less frequently displaced, leading to a more static hierarchy of engagement.
In sum, user location disclosure reconfigured both the meaning and the risks of online discourse. By making regional identity visible, it raised the cost of participation—especially across provincial boundaries—resulting in lower comment volume, diminished dissent, and a narrower, more fragmented public conversation.
4.4 Robustness checks
@@
We conduct a series of robustness checks, detailed in the SI, to verify the stability of our main findings on the decline in comment volume due to the implementation of the user location disclosure policy. First, we vary the bandwidths used in the local smoothing, including those selected by the rdrobust package, bandwidths chosen via cross-validation, and a range of fixed windows (8, 24, 72, 96, and 120 hours). In all cases, the observed decline in comment volume remains consistent. Second, we re-estimate the models using a log-transformed outcome, $\log(\#comments+1)$, to assess sensitivity to outliers. While the main analyses use the number of unique comments to accommodate posts with zero comments, the log transformation yields qualitatively similar results. Finally, we perform placebo tests by shifting the timing of the policy intervention to alternative dates. These placebo specifications show no discontinuities, reinforcing that the observed effects are indeed tied to the actual timing of the location disclosure policy.
5 Conclusion
We study the impact of a user location disclosure policy on Sina Weibo, focusing on how the introduction of IP-based location tags affected online engagement. Although framed as a way to curb misinformation and foreign influence, the policy had little effect on overseas participation. Instead, it sharply reduced domestic users’ critical engagement with out-of-province issues. It appears that this decline was driven not by fear of state repression but by increased exposure to regionally discriminatory backlash.
These findings show that authoritarian regimes can shape discourse not only through direct censorship but also by enabling peer-based sanctions via platform design. Making identity visible can unintentionally activate social cleavages that suppress participation and narrow the scope of expression. While this may help stabilize discourse in the short run, it risks muting early signals of discontent, with potential costs for the quality of governance (Chen and Xu, 2017b).
More broadly, our study contributes to a growing literature on how platform architecture shapes user behavior. Research on algorithmic curation, identity labeling, and credibility signals shows that seemingly minor design choices can affect exposure, trust, and sharing (Guess et al., 2023; Liang et al., 2022; Nassetta and Gross, 2020; Papakyriakopoulos et al., 2022; Bradshaw et al., 2024). We extend this work by demonstrating that identity cues alone can alter participation even without content removal or fact correction. This suggests that platform design decisions—whether in democracies or authoritarian states—can shape the boundaries of public discourse in subtle, unintended, but consequential ways.
@@ 5 Conclusion
We monitored the timelines of over 200 official Weibo accounts in real time, spanning various levels of Chinese government bodies and media outlets, as illustrated in Panel 1 of the main text. Table LABEL:table:account_list lists the 165 accounts included in this study, along with their corresponding English names.
@@
Notes: Panel 4(a) shows the average number of new reads for a Weibo post tracked continuously for 24 hours after publication. Readership peaks immediately after posting, with an average of about 12,000 new reads, then drops sharply within the first four hours and stabilizes at a much lower level. This pattern is consistent across our full dataset of 94,009 posts. Panel 4(b) presents the corresponding trend in comment activity. Comments also peak shortly after publication, though at a smaller scale, averaging about two at the peak, and decline rapidly to near zero within five hours. These patterns validate our monitoring strategy—five-minute intervals during the first 24 hours and daily snapshots thereafter—as sufficiently granular to capture the vast majority of user engagement. Panel 4(c) and Panel 4(d) display the hourly distribution of post and comment snapshots aggregated across the observation period, with activity peaking between 08:00 and 12:00—typical morning posting hours. shows the hourly distribution of comment snapshots, with most activity occurring between 06:00 and 22:00, reflecting expected user engagement hours.
@@
Notes: This figure illustrates the model used to classify the sentiment of events described in Weibo posts from the perspective of the Chinese government. Posts were processed using a structured prompt instructing the model to label each event as "positive," "neutral," or "negative" with no explanatory text. The model used was GPT-4-turbo-preview, with fixed parameters to ensure deterministic outputs. The example shown involves a Weibo post reporting a fatal accident, which the model correctly classifies as "negative." This approach allows for consistent and scalable sentiment classification across a large corpus of posts. Translation of Prompt: "Please analyze the following Weibo content and determine, from the perspective of the Chinese government, whether the event described is ’positive,’ ’neutral,’ or ’negative.’ Output only one of these three categories without additional text. Here is the content of the Weibo post: {post}." Translation of Info: "[#3 dead, 5 injured as SUV plunges into park lake in Tianjin#] According to Qingjiao Video Report: Around 15:00 on April 16, according to CCTV news #Beichen Yedi Park (Yongding Xinhe Riverside)#, an accident occurred. An SUV with five people on board accidentally slid into the river embankment and collided with three citizens resting by the riverbank, then fell into the Yongding Xinhe. The accident caused 3 deaths on-site, and 5 injured were sent to the hospital for treatment." Translation of Output: "Negative".
@@
[…]
@@
[…]
@@
Taking the square root of the appropriate diagonal element yields the cluster-robust standard error $\hat{\tau}$, $\widehat{\mathrm{se}}(\hat{\tau})$. The confidence intervals are obtained through normal approximation.
@@
For a candidate bandwidth $h$ we perform $K$-fold block cross-validation: split the running variable into $K$ non-overlapping bins that are outside an exclusion zone around $t_{0}$, re-estimate Model (1) on $K-1$ bins, and compute the root-mean-squared prediction error (RMSE) on the hold-out bin. We then choose
@@
Table S7 presents the estimated causal effects of user location disclosure on a range of outcomes across multiple bandwidths. These results extend the findings shown in Figures 2 and 3 of the main text. The CV Bandwidth” column reports the optimal bandwidth selected using the rdrobust() function from the R package rdrobust. The RD Bandwidth” and “RD” columns show the corresponding bandwidths and treatment effect estimates. To assess robustness, we also report estimates using alternative bandwidths of 8, 24, 72, 96, and 120 hours. The estimated effects are consistent across specifications, which shows that our main estimates are robust.
@@
Table S7: Treatment Effect Results Table
@@
Figure 2(a) Topic: International #Comments 323232 0.369 107 0.342 0.545 0.407 0.396 0.435 0.454
@@
Figure 2(a) Topic: Non-International #Comments 323232 -0.011 93 -0.010 0.214 0.005 0.004 0.019 0.027
@@
Figure 2(b) Topic: International #Comments 969696 0.407 122 -2.078 0.378 -1.783 -0.191 0.407 0.661
@@
Figure 2(b) Topic: Non-International #Comments 161616 -4.110 95 -4.603 -0.019 -4.607 -3.711 -3.353 -3.178
@@
Figure 2(c) Topic: Non-Local #Comments 404040 -1.399 89 -1.473 1.201 -1.511 -0.631 -0.194 0.030
@@
Figure 2(c) Topic: Local #Comments 323232 -7.794 99 -7.793 -0.887 -7.599 -6.748 -6.392 -6.224
@@
Figure 2(e) User: In-Province #Comments 484848 -1.710 88 -1.896 -0.121 -1.869 -1.500 -1.426 -1.399
@@
Figure 2(e) User: Out-of-Province #Comments 404040 -5.665 103 -5.702 -0.801 -5.559 -5.008 -4.735 -4.600
@@
Figure 2(e) User: Oversea #Comments 969696 -0.073 96 -0.108 0.060 -0.100 -0.086 -0.073 -0.066
@@
Figure 2(f) User: In-Province #Critical Comments 404040 -0.572 88 -0.615 -0.436 -0.596 -0.410 -0.370 -0.355
@@
Figure 2(f) User: Out-of-Province #Critical Comments $8$3.591 38 -8.153 -7.879 -6.730 -4.181 -3.430 -3.195
@@
Figure 2(f) User: Oversea #Critical Comments 969696 -0.030 81 0.004 0.026 -0.001 -0.036 -0.030 -0.025

### v3

@@ 4.1 Data and Measurement
We conducted real-time monitoring of each account’s timeline (i.e., their posts), each post’s first-page comment section (displaying up to the top 20 comments), and associated engagement metrics at both the post and comment levels. We focus on the first-page comments because they were instantly available to readers when a post is clicked and draw the most user attention and interaction. As shown in Figure 1(a), our system scanned all monitored accounts, as well as recorded posts and comments, every five minutes. This allowed us to collect longitudinal data at both the account and post levels. When a new post appeared, the system initiated tracking and began recording data on the post and its top comments. Snapshots were captured every five minutes for the first 24 hours, and once every day for the following ten days. While most engagement occurred within the first few hours (see Figure B3 in the Supplementary Information), the extended tracking window enabled us to capture delayed or atypical activity. Figure 1(b) illustrates the variables collected through scraping, including each post’s content, publication time, and engagement metrics (number of reposts, likes, and comments), as well as the first-page of the comment section (top 20 comments). For each comment and their replies, we also recorded the associated IP location tag. Figure 1(c) and Figure 1(d) show the numbers of new posts, and new comments collected within a 48-hour window, respectively. Figures 1(e) and 1(f) show that, on average, comments appearing on and after the second page receive minimal engagement.
Timeline and Location Tags
@@
The location tags reflect users’ IP addresses at the time of commenting, although we do not observe or record the IPs themselves. Overseas users are labeled by country or region, while domestic users are tagged at the provincial level (e.g., Shanghai, Jiangsu, Shaanxi). [^3] A potential concern is that users may mask their IP addresses using VPNs. in the Supplementary Information, we show that fewer than 1% of users changed location tags during the study period. VPN usage among domestic users is rare and typically serves to switch between domestic and foreign IPs, not between provinces. To our knowledge, commercially available VPN services in China do not support intra-national IP masking, as all provinces operate within the same national firewall. These patterns suggest that provincial-level location tags are generally reliable. Excluding users whose location tags have changed does not alter on our findings.
@@
Post content: positive, negative, or neutral event Post LLM-assisted annotation + human validation $0.721$
@@
Comment stance: critical of the post or not Comment Supervised BERT classifier + human validation $0.783$
Comment receiving regionally discriminatory replies Comment LLM-assisted annotation + human validation $0.835$
@@
We also classify post types based on their content. International posts are those that mention one of 123 countries or regions outside China. We construct the list of 123 countries and regions (see Table D4 in the Supplementary Information) by extracting all foreign country and region names that appeared in Weibo’s post-treatment IP location tags. Local posts reference one of China’s 31 provinces. We focus on provincial references, rather than prefectures or counties, because IP location tags are reported at the country level for overseas users and the provincial level for domestic users. Posts are further categorized as depicting a positive, negative, or neutral event (from the perspective of the Chinese government) using a large language model (LLM), followed by human validation.
@@
To detect regional discrimination in replies, we use LLM-based annotation, validated by human coders, to identify whether a reply contains discriminatory content based on the commenter’s location. We also track the frequency with which a user’s IP location is mentioned in replies as an indirect measure of the salience of geographic identity in the discourse. Table 1 summarizes these variables and their corresponding measurement strategies. Table C3 in the Supplementary Information provides the summary statistics.
Protecting Privacy
All data used in this study are publicly available. Nevertheless, we take additional steps to protect the privacy of commenters, many of whom are likely ordinary citizens in China. First, we do not collect any personally identifiable information, such as commenters’ IP addresses. The location tags we use are limited to the provincial level and are too coarse to identify individuals. Second, we mask all user identifiers—including usernames, avatars, and URLs—by replacing them with randomly assigned IDs to ensure complete anonymity. Third, all data are stored on a secure server at the corresponding author’s institution, with access restricted to authorized personnel only (see Table B2 in the Supplementary Information for details.)
4.2 Causal Identification
@@
Because an interrupted time series design is analogous to a regression discontinuity design with time as the forcing variable (Hausman and Rapson, 2018), we estimate $\tau_{\text{SAITT}}$ using local linear regression on either side of $c$, clustering standard errors at the account level. Bandwidths are selected using both the MSE-optimal procedure in rdrobust (Calonico et al., 2014) and cross-validation. We show in the Supplementary Information Section F that our results are robust to a wide range of bandwidth choices.
Robustness checks
@@
We conduct a series of robustness checks, detailed in the Supplementary Information, to verify the stability of our main findings on the decline in comment volume due to the implementation of the user location disclosure policy. First, we vary the bandwidths used in the local smoothing, including those selected by the rdrobust package, bandwidths chosen via cross-validation, and a range of fixed windows (8, 24, 72, 96, and 120 hours). In all cases, the observed decline in comment volume remains consistent. Second, we re-estimate the models using a log-transformed outcome, $\log(\#comments+1)$, to assess sensitivity to outliers. While the main analyses use the number of unique comments to accommodate posts with zero comments, the log transformation yields qualitatively similar results. Finally, we perform placebo tests by shifting the timing of the policy intervention to alternative dates. These placebo specifications show no discontinuities, reinforcing that the observed effects are indeed tied to the actual timing of the location disclosure policy.
@@
We monitored the timelines of over 200 official Weibo accounts in real time, spanning various levels of Chinese government bodies and media outlets. Table LABEL:table:account_list lists the 165 accounts included in this study, along with their corresponding English names.
@@
Notes: Panel 4(a) shows the average number of new reads for a Weibo post tracked continuously for 24 hours after publication. Readership peaks immediately after posting, with an average of about 12,000 new reads, then drops sharply within the first four hours and stabilizes at a much lower level. This pattern is consistent across our full dataset of 94,009 posts. Panel 4(b) presents the corresponding trend in comment activity. Comments also peak shortly after publication, though at a smaller scale, averaging about two at the peak, and decline rapidly to near zero within five hours. These patterns validate our monitoring strategy—five-minute intervals during the first 24 hours and daily snapshots thereafter—as sufficiently granular to capture the vast majority of user engagement. Panel 4(c) and Panel 4(d) display the hourly distribution of post and comment snapshots aggregated across the observation period, with activity peaking between 08:$0$ and 12:00—typical morning posting hours. shows the hourly distribution of comment snapshots, with most activity occurring between 06:$0$ and 22:$0$, reflecting expected user engagement hours.
@@
Notes: This figure illustrates the model used to classify the sentiment of events described in Weibo posts from the perspective of the Chinese government. Posts were processed using a structured prompt instructing the model to label each event as "positive," "neutral," or "negative" with no explanatory text. The model used was GPT-4-turbo-preview, with fixed parameters to ensure deterministic outputs. The example shown involves a Weibo post reporting a fatal accident, which the model correctly classifies as "negative." This approach allows for consistent and scalable sentiment classification across a large corpus of posts. Translation of Prompt: "Please analyze the following Weibo content and determine, from the perspective of the Chinese government, whether the event described is ’positive,’ ’neutral,’ or ’negative.’ Output only one of these three categories without additional text. Here is the content of the Weibo post: {post}." Translation of Info: "[#3 dead, 5 injured as SUV plunges into park lake in Tianjin#] According to Qingjiao Video Report: Around 15:$0$ on April 16, according to CCTV news #Beichen Yedi Park (Yongding Xinhe Riverside)#, an accident occurred. An SUV with five people on board accidentally slid into the river embankment and collided with three citizens resting by the riverbank, then fell into the Yongding Xinhe. The accident caused 3 deaths on-site, and 5 injured were sent to the hospital for treatment." Translation of Output: "Negative".
@@
Let $y_{gi}$ denote the outcome for observation $i$ in cluster (account) $g$, and let $t_{gi}$ be the corresponding timestamp measured in seconds. We are interested in estimating the instantaneous treatment effect at the cutoff $t_{0}=\text{2022-04-28 12{:}00}$. Define the centered running variable as $\Delta_{gi}=t_{gi}-t_{0}$, and let the treatment indicator be $D_{gi}=\mathbbm{1}{\Delta_{gi}\geq 0}$.
@@
where $\gamma_{g}$ captures any time-invariant heterogeneity at the account level. Model 1 is estimated only with observations whose $|\Delta_{gi}|$ lies in a bandwidth window [h,h][-h,h]; observations are moreover re-weighted by a Gaussian kernel:
@@
[…]
@@
For a candidate bandwidth $h$ we perform $K$-fold block cross-validation: split the running variable into $K$ non-overlapping bins that are outside an exclusion zone around $t_{0}$, re-estimate Model (1) on K1K-1 bins, and compute the root-mean-squared prediction error (RMSE) on the hold-out bin. We then choose
@@
Table S7 presents the estimated causal effects of user location disclosure on a range of outcomes across multiple bandwidths. These results extend the findings shown in Figures 1 and 2 of the main text. The CV Bandwidth” column reports the optimal bandwidth selected using the rdrobust() function from the R package rdrobust. The RD Bandwidth” and “RD” columns show the corresponding bandwidths and treatment effect estimates. To assess robustness, we also report estimates using alternative bandwidths of 8, 24, 72, 96, and 120 hours. The estimated effects are consistent across specifications, which shows that our main estimates are robust.
@@
Table S7: Treatment Effect Estimates
@@
Figure 1a Topic: International #Comments $32$ 0.369 107 0.342 0.545 0.407 0.396 0.435 0.454
@@
Figure 1a Topic: Non-International #Comments $32$ -0.011 93 -0.010 0.214 0.005 0.004 0.019 0.027
@@
Figure 1b Topic: International #Comments $96$ 0.407 122 -2.078 0.378 -1.783 -0.191 0.407 0.661
@@
Figure 1b Topic: Non-International #Comments $16$ -4.110 95 -4.603 -0.019 -4.607 -3.711 -3.353 -3.178
@@
Figure 1c Topic: Non-Local #Comments $40$ -1.399 89 -1.473 1.201 -1.511 -0.631 -0.194 0.030
@@
Figure 1c Topic: Local #Comments $32$ -7.794 99 -7.793 -0.887 -7.599 -6.748 -6.392 -6.224
@@
Figure 1e User: In-Province #Comments $48$ -1.710 88 -1.896 -0.121 -1.869 -1.500 -1.426 -1.399
@@
Figure 1e User: Out-of-Province #Comments $40$ -5.665 103 -5.702 -0.801 -5.559 -5.008 -4.735 -4.600
@@
Figure 1e User: Oversea #Comments $96$ -0.073 96 -0.108 0.060 -0.100 -0.086 -0.073 -0.066
@@
Figure 1f User: In-Province #Critical Comments $40$ -0.572 88 -0.615 -0.436 -0.596 -0.410 -0.370 -0.355
@@
Figure 1f User: Out-of-Province #Critical Comments $88$ -3.591 38 -8.153 -7.879 -6.730 -4.181 -3.430 -3.195
@@
Figure 1f User: Oversea #Critical Comments $96$ -0.030 81 0.004 0.026 -0.001 -0.036 -0.030 -0.025